solve each equation by completing the square b^2 -12b +45 = 10
Answer by waynest(281)

b^2 - 12b + 45 = 10
b^2 - 12b + 45 - 10 = 10 - 10
b^2 - 12b + 35 = 0
b^2 - 12b + (12/2)^2 - (12/2)^2 + 35 = 0
b^2 - 12b + 36 - 36 + 35 = 0
(b - 6)^2 - 1 = 0
(b - 6)^2 = 1
sqrt(b - 6)^2 = +-sqrt1
b - 6 = +- 1
b - 6 + 6 = 6 +- 1
b = 6+ 1
b = 7
b = 6 - 1
b = 5